# On the distance to the black hole X-ray binary Swift J1727.8--1613

### v1

@@ 4.3 Updated $E(B\!-\!V)$ constraints and resulting distance
To further constrain the value of $E(B\!-\!V)$, we first use Galactic dust maps to compute the change in $E(B\!-\!V)$ along the line of sight to J1727. We present these results in Figure 4, alongside our constraint of $E(B\!-\!V)=0.37\pm 0.01\,({stat})\pm 0.025\,({sys})$ derived from near-UV observations (Section 3.2). Our results imply that $E(B\!-\!V)$ cannot be much greater than 0.4, as we run out of dust along the line of sight.
@@
Figure 4: Estimates of $E(B\!-\!V)$ for distances between 007.57.57.57.5 kpc along the line of sight to J1727. The 3D Galactic dust maps include Bayestar (blue line; Green et al., 2019) and Gaia (purple line; Edenhofer et al., 2024). We also use an update of the 2D “SFD” dust map (Schlegel et al., 1998) to remove extragalactic contamination that is referred to as the “corrected SFD” (CSFD; grey dashed line; Chiang, 2023). While Green et al. (2019) and Edenhofer et al. (2024) have significant caveats, a comparison between the two clearly shows a more rapid climb in $E(B\!-\!V)$ at closer distances for the Gaia-based map. To investigate the $E(B\!-\!V)$ values that can be reached at $\sim 5$ kpc, we use CSFD, which only reports the total value for $E(B\!-\!V)$ along the line of sight, and therefore serves as a good estimate of the maximum $E(B\!-\!V)$ in the direction of J1727 at high distances. From this we derive a maximum $E(B\!-\!V)\approx 0.4$. Despite disagreement between Bayestar and Gaia in how gas is distributed along the line of sight, both appear to accumulate to roughly this maximum value. Lastly, our estimate of $E(B\!-\!V)$ (pink dotted line) is derived from our near-UV observations. We use this estimate to conclude that values of $E(B\!-\!V)\gg 0.4$ may not be reliable, and suggest a distance to J1727 of $5.5^{+1.4}_{-1.1}$ kpc, which is included for reference (orange dot-dash line).
We note that MS24b calculate a best estimate of the distance to J1727 by assuming equal weights for all methods. Given the unlikely applicability of method 1, and the implausibility of the $E(B\!-\!V)$ values obtained with methods 2 and 3, we suggest using an updated version of method 4.
The reported value for $N_{\rm{H}}$ (Draghis et al., 2023) has been estimated using the tbabs X-ray absorption model and Wilms et al. (2000) abundances of elements. Therefore, the correlation reported by Zhu et al. (2017) becomes a more appropriate choice for estimating extinction than Güver & Özel (2009). Assuming $R_{V}=3.1$ and using equation 11 from Zhu et al. (2017), we estimate $E(B\!-\!V)=0.36\pm 0.11$ along the line of sight to J1727. This is in remarkable agreement with our findings based on near-UV data.
Resampling parameters $10^{5}$ times using Monte Carlo techniques, and using Equation 1 with the MS24b values for $M_{r}$ and $m_{r}$, our estimate of $E(B\!-\!V)$ from $N_{\rm{H}}$ gives a median distance of $5.5^{+1.6}_{-1.3}$ kpc, as compared to a value of $5.5^{+1.4}_{-1.1}$ kpc from our more precise near-UV-derived value. These distance estimates correspond to a GP elevation of $z\approx 1.0^{+0.3}_{-0.2}$ kpc, and a high likelihood that the H i density near J1727 will be well below the level required for H i absorption to be detectable. We collate these distances alongside other estimates in Table 2.
Reference Method Reported Distance (kpc)
Abdulghani et al. (2024) Eddington luminosity fraction $1.52^{+0.85}_{-0.61}$
Mata Sánchez et al. (2024) Interstellar Ca ii doublet equivalent width ratio $3.2\pm 0.6$
Mata Sánchez et al. (2025) $E(B\!-\!V)$ from interstellar K i 7699 Å  equivalent width $3.2\pm 1.3$
Mata Sánchez et al. (2025) $E(B\!-\!V)$ from diffuse interstellar band 8621 Å  equivalent width $3.5\pm 1.4$
Mata Sánchez et al. (2025) $E(B\!-\!V)$ from $N_{\rm{H}}$ per Güver & Özel (2009) $4.8\pm 1.4$
This work H i absorption kinematic distance $3.6\pm 0.3\,({stat})\pm 2.3\,({sys})$
This work $E(B\!-\!V)$ from $N_{\rm{H}}$ per Zhu et al. (2017) $5.5^{+1.6}_{-1.3}$
This work $\boldsymbol{E(B\!-\!V)}$ from HST/STIS near-UV observations $\mathbf{5.5^{+1.4}_{-1.1}}$
Table 2: A comparison of J1727 distance estimates arising from different methods. The ELF-derived distance is likely underestimated (Abdulghani et al., 2024). Where MS24a estimates have been updated since, we use the MS24b values. These both report mean values and are further discussed in Sec. 4.2. The Ca ii method is unlikely to meet the condition that the J1727 is less than a few kpc away from the GP as shown in Sections 4.1.2 and 4.3. The equivalent widths at 7699 Å  and 8621 Å  are both derived using relations from scattered data and infer values of $E(B\!-\!V)$ that are high enough to be unlikely. MS24a acquires an estimate of $N_{\rm{H}}$ from X-ray studies of J1727 and uses $\langle N_{\rm{H}}/A_{V}\rangle$ from Güver & Özel (2009) to calculate $E(B\!-\!V)$, which is reused by MS24b to calculate the distance. The H i absorption kinematic distance derived in this work is a lower bound on the distance as there could be additional distance between the furthest observed H i cloud and J1727. This estimate accounts for the impact of Galactic longitude on kinematic distance methodologies in the form of additional systematic uncertainty. We use the more recent results of Zhu et al. (2017) to implement a larger value of $\langle N_{\rm{H}}/A_{V}\rangle$ and recalculate the distance implied by $N_{\rm{H}}$. Lastly, we use near-UV observations to constrain $E(B\!-\!V)$ and subsequently the distance shown in bold as our preferred estimate. The distance estimates from this work are median values with $1\sigma$ uncertainties.
4.4 Distance implications
Wood et al. (2024) resolved a jet of size $\sim\!110(d/2.7\,\rm{kpc})/\rm{sin}\,\mathit{i}$ AU at 8.4 GHz, where $d$ is the distance and $i$ is the jet axis inclination angle. Our suggested distance of $5.5^{+1.4}_{-1.1}$ kpc suggests a jet size $>200$ AU regardless of the inclination angle used.
Liu et al. (2024) found that in the hard state at the beginning of the J1727 outburst, the broadband X-ray flux was $\sim 3\times 10^{-7}\ \rm{erg\ cm^{-2}\ s^{-1}}$. At a distance of $5.5^{+1.4}_{-1.1}$ kpc, this would correspond to a luminosity of $\sim 1.1\times 10^{39}\ \rm{erg\ s^{-1}}$, which is equivalent to $L_{\rm{Edd}}$ for an $8.6\,\rm{M}_{\odot}$ BH accreting hydrogen. While these luminosities are unusually high for the pure hard state, it is not impossible (Tetarenko et al., 2016).
The inferred natal kick velocity of J1727 is also impacted by an increase in the estimated distance. MS24b used their estimated distance of $3.4\pm 0.3$ kpc, along with the proper motion of J1727, to derive a median and 68% confidence level in the potential natal kick velocity of $v_{\rm{kick}}=210^{+40}_{-50}$ km s-1. Our new distance revises this to $v_{\rm{kick}}=190\pm 30$ km s-1(Atri et al., 2019) [^7] https://github.com/pikkyatri/BHnatalkicks.
@@ 5 Conclusions
Using H i absorption data from MeerKAT observations of the outburst of Swift J1727.8--1613, we determine a maximum absorption velocity of $24.8\pm 2.8$ km s-1. The higher-velocity absorption seen towards the extragalactic reference source PKS J1733--1304 allows us to use the near kinematic distance as a lower bound of $3.6\pm 0.3\,({stat})\pm 2.3\,({sys})$ kpc, accounting for the systematic uncertainty due to the low Galactic longitude of the source. Its high Galactic latitude likely implies that we run out of H i clouds along the line of sight, suggesting that the true distance could be significantly greater.
We therefore make use of a near-UV spectrum of Swift J1727.8--1613 as observed by the Space Telescope Imaging Spectrograph aboard the Hubble Space Telescope. This allows us to estimate the colour excess or reddening as $E(B\!-\!V)=0.37\pm 0.01\,({stat})\pm 0.025\,({sys})$. This is significantly lower than previous constraints on $E(B\!-\!V)$ from Mata Sánchez et al. (2024, 2025), but is in good agreement the maximum value derived from Galactic dust maps along this line of sight, and suggests that previous works underestimated the distance.
Using $E(B\!-\!V)$ to determine the $r$-band extinction and combining this with donor star $r$-band magnitudes measured by Mata Sánchez et al. (2025), we present $5.5^{+1.4}_{-1.1}$ kpc as the resulting and likely distance to Swift J1727.8--1613. This places the system much farther away than previously thought and implies a natal kick velocity of $v_{\rm{kick}}=190\pm 30$ km s-1.
@@ Acknowledgements
We thank James Allison for his development of the ThunderKAT H i Pipeline.
The MeerKAT telescope is operated by the South African Radio Astronomy Observatory, which is a facility of the National Research Foundation, an agency of the Department of Science and Innovation.
@@
This work made use of the CARTA (Cube Analysis and Rendering Tool for Astronomy) software (DOI 10.5281/zenodo.3377984https://cartavis.github.io).

### v2

@@ 4.2.1 E(BV)E(B\!-\!V) validation
To first validate our result of $E(B\!-\!V)=0.37\pm 0.01\,({stat})\pm 0.025\,({sys})$ as derived from near-UV observations in Section 3.2, we compute the change in E(BV)E(B\!-\!V) along the line of sight to J1727 using Galactic dust maps. We present these results in Figure 4. Our results imply that E(BV)E(B\!-\!V) cannot be much greater than 0.4, as we run out of dust along the line of sight (but note the spatial resolution of dust maps could be a limiting factor). In fact, our best value for E(BV)E(B\!-\!V) is lower than all those estimated by MS24.
@@
Figure 4: Estimates of E(BV)E(B\!-\!V) for distances between $0\text{--}7.5$ kpc along the line of sight to J1727. The 3D Galactic dust maps include Bayestar (blue line; Green et al., 2019) and Gaia (purple line; Edenhofer et al., 2024). We also use an update of the 2D “SFD” dust map (Schlegel et al., 1998) to remove extragalactic contamination that is referred to as the “corrected SFD” (CSFD; grey dashed line; Chiang, 2023). To investigate the likely maximum E(BV)E(B\!-\!V) values within this distance range, we use CSFD; as this only reports the total value for E(BV)E(B\!-\!V) along the line of sight, it serves as a good estimate of the maximum E(BV)E(B\!-\!V) in the direction of J1727 at high distances. From this we derive a maximum $E(B\!-\!V)\approx 0.4$. Despite each having significant caveats and the disagreement between Bayestar and Gaia in how gas is distributed along the line of sight, both appear to accumulate to roughly this maximum value. Lastly, our estimate of E(BV)E(B\!-\!V) (pink dotted line) is derived from our near-UV observations. We use this estimate to conclude that values of $E(B\!-\!V)\gg 0.4$ may not be reliable, and suggest a distance to J1727 of $5.5^{+1.4}_{-1.1}$ kpc with $1\sigma$ uncertainties, which are included for reference (orange dot-dash line and shaded region respectively).
We note that while there is no evidence for intrinsic variability in $N_{\rm H}$, the values used in studies of J1727 have varied from as low as $1.2\pm 0.2\times 10^{21}\,{\rm cm^{-2}}$ (Chatterjee et al., 2024) to as high as $(4.1\pm 0.1)\times 10^{21}\,{\rm cm^{-2}}$ (Draghis et al., 2023) with other values in between (O’Connor et al., 2023; Peng et al., 2024; Svoboda et al., 2024). These discrepancies are likely due to the instrumental systematics and/or modelling choices differing from study to study.
MS24 used $(3.2\pm 0.9)\times 10^{21}\,{\rm cm^{-2}}$, derived as the mean $N_{\rm H}$ value from O’Connor et al. (2023) and Draghis et al. (2023). The latter was estimated using the tbabs X-ray absorption model and Wilms et al. (2000) abundances of elements. Both were derived by fitting observations made in late August 2023, when the source was in the rising hard/hard-intermediate state.
We instead use $(2.4\pm 0.1)\times 10^{21}\,{\rm cm^{-2}}$ per Svoboda et al. (2024), which was derived using NICER, NuSTAR, and IXPE observations made during the soft state. The broad spectral coverage of this data set and the well-characterised disk blackbody spectrum during the soft state allows for a more robust estimate of $N_{\rm H}$ than those derived from a more restricted bandpass or made during the hard or intermediate states. This value is in agreement with the HI4PI Survey’s measured $N_{\rm H}$ in the direction of J1727 (HI4PI Collaboration et al., 2016), and reasonable agreement with the $N_{\rm H}$ values presented by Chatterjee et al. (2024) and Peng et al. (2024).
We also use an alternative relation, $N_{\rm H}({\rm cm}^{-2})=(2.87\pm 0.04)\times 10^{21}A_{V}{\rm\,(mag)}$ (Zhu et al., 2017, Equation 11), rather than that of Güver & Özel (2009). This more recent relation is a more appropriate choice for estimating extinction, given it was determined by fitting the whole sample of Wilms et al. (2000) abundances, which include XRBs. However, we note there is likely additional uncertainty due to the scatter observed in this relation (Zhu et al., 2017, Figure 9(a)). Assuming $R_{V}=3.1$, we use Monte Carlo techniques and Equation 2 to instead estimate $E(B\!-\!V)=0.27\,\pm\,0.01$ along the line of sight to J1727. This result, while dependent upon the choice of $N_{\rm H}$ and with likely greater uncertainty, are in agreement with our findings based on near-UV data and Galactic dust maps that $E(B\!-\!V)\lesssim 0.4$.
4.2.2 Suggested distance to Swift J1727.8-1613
With $A_{r}=0.84\pm 0.06$ per Section 3.2, the remaining parameter values required for Equation 1 are the absolute and apparent $r$-band magnitudes, $M_{r}$ and $m_{r}$ respectively. MS25 determine the donor star type to be K4($\pm$1)V with $M_{r}=6.6\pm 0.5$ mag and $m_{r}=21.13\pm 0.05$ mag. Using these values produces the distance posterior distribution that we present in Figure 5 and a distance estimate of $5.5^{+1.4}_{-1.1}$ kpc, under the assumption that the donor star is a regular K4($\pm$1)V star that has not lost significant mass during the binary evolution. In such a case, the distance would be reduced, implying reduced values of E(BV)E(B\!-\!V) and $N_{\rm H}$.
4.2.3 Nature of the donor star
Using our distance estimate and the results from Wood et al. (2025), we define an upper limit on the inclination angle, $i$, for J1727 of $i\lesssim 69^{\circ}$. To define a conservative lower limit on $i$, we assume an upper limit of $20\,{\rm M_{\odot}}$ for the black hole mass. It is likely that the black hole is actually much smaller, as discussed with reference to high natal kick velocity in Section 4.3. We also use a maximum-entropy skew-normal distribution for the K4($\pm$1)V donor star mass such that $1\sigma$ is contained within the range $0.70\text{--}0.78\,{\rm M_{\odot}}$ and $\mu\approx 0.73\,{\rm M_{\odot}}$. These mass values allow us to use the mass function provided by MS25,
$f(M)\equiv\frac{M_{1}\sin^{3}i}{(1+M_{2}/M_{1})^{2}}=2.77\pm 0.09{\rm M_{\odot}},$ (3)
to estimate a $1\sigma$ lower limit of $i\geq 32.0\,\pm\,0.4^{\circ}$.
MS25 obtained an upper limit of $v\sin i<52$ km s-1 for the rotational broadening of the donor star, but recommend a more conservative $3\sigma$ upper limit of $102$ km s-1 as this constraint is not especially strong. Values of $v\sin i>102$ km s-1 are therefore unlikely, and would require $M_{2}>1M_{\odot}$.
Given the binary system’s orbital period, $P_{\rm orb}$, and assuming that the donor star is tidally locked, the donor star’s radius can be calculated as
$R_{2}=\frac{P_{\rm orb}(v\sin i)}{2\pi\sin i}.$ (4)
MS25 directly measured $P_{\rm orb}=10.8038\pm 0.0010$ hrs. Combining this with the $v\sin i=27^{+25}_{-21}$ km s-1 posterior distribution from MS25 and a uniform distribution of $\cos i\sim U(\cos 69^{\circ},\cos 32^{\circ})$ results in an $R_{2}$ distribution where $R_{2}\ll 1.0\,{\rm R_{\odot}}$, whereas K4($\pm$1)V stars typically have radii of $R\approx 1\,{\rm R_{\odot}}$.
Paczyński (1967) relates the XRB orbital period and the mass and radius of a donor star undergoing Roche lobe overflow as
$R_{L}\approx(2GM_{2})^{1/3}(P_{\rm orb}/9\pi)^{2/3}$ (5)
where $R_{L}$ is the Roche lobe radius, and $M_{2}$ is the donor star (secondary) mass. Taking $R_{L}=R_{2}$ as derived using the $v\sin i$ posterior distribution from MS25 results in an $M_{2}$ distribution that also favours extremely small, unphysical values ($M_{2}\ll 0.1\,{\rm M_{\odot}}$).
Donor stars of BH low-mass XRBs may be significantly evolved (Podsiadlowski et al., 2003; Fragos & McClintock, 2015) and in Roche lobe overflow. However, there is no current evidence from the donor star spectral analysis by MS25 that it is significantly evolved or stripped, which would impact the distance estimate. Only with future, higher resolution spectroscopic observations could we gain further information into the nature of the J1727 donor star.
We therefore infer that $v\sin i$ is unlikely to be as low as the MS25 posterior distribution might suggest. Assuming co-rotation and Roche lobe overflow, we again combine equations 4 and 5. Then, using $P_{\rm orb}$ from MS25 and our aforementioned $M_{2}$ skew-normal and $\cos i$ uniform distributions, we calculate $v\sin i$ to be in the $3\sigma$ range $60\text{--}112$ km s-1. This is roughly consistent with MS25 $3\sigma$ upper limit $v\sin i\lesssim 102$ km s-1, given it is dependent upon conservative limits on $i$, and further contradicts the abundance of low values in the MS25 $v\sin i$ posterior.
Figure 5: The distance posterior distribution produced from Monte Carlo calculations using the near-UV constraint on E(BV)E(B\!-\!V). The solid vertical line represents the 0.5000 quantile and the dashed vertical lines represent the 0.1585 and 0.8415 quantiles.
4.3 Implications
4.3.1 Natal kick velocity and birth mechanism
Using their estimated distance of $3.4\pm 0.3$ kpc, along with the proper motion of J1727, MS25 derived a median and $1\sigma$ confidence level in the potential natal kick velocity of $v_{\rm{kick}}=210^{+40}_{-50}$ km s-1. We use our suggested distance of $5.5^{+1.4}_{-1.1}$ kpc to provide an updated natal kick velocity of $190\,\pm\,30$ km s-1 (Atri et al., 2019) [^7] https://github.com/pikkyatri/BHnatalkicks. Our revised distance also corresponds to Galactic elevations of $z\approx 0.8\text{--}1.2$ kpc. We can compare this to XTE J1118+480, for which Gualandris et al. (2005) presented an asymmetric natal kick with a similar velocity of $183\pm 31$ km s-1 and a Galactic elevation of $1.9\pm 0.4$ kpc. Atri et al. (2019) estimated a similar natal kick velocity for XTE J1118+480 and particularly high velocity kicks for other systems, such as 4U 1543-475, GS 1354-64, and SAX J1819.3-$25$, and suggested that these are indicative of asymmetrical supernovae explosions as the likely birth mechanism. Applying this to J1727, it is possible that it could have been born within the Galactic disk, and propelled by a natal kick to its present-day Galactic elevation. It is also possible that the large natal kick has caused a misalignment between the BH spin and the accretion disk for J1727 (Maccarone, 2002; Martin et al., 2008). However, no evidence has been observed to date of precession of the jet axis (Wood et al., 2024, 2025).
4.3.2 Eddington luminosity fraction
Zdziarski et al. (2025) calculate an unabsorbed bolometric flux of $F_{\rm bol}\approx 5\times 10^{-7}\,{\rm\,erg\,cm^{-2}\,s^{-1}}$ for J1727, as informed by the peak hard state flux from Liu et al. (2024). At a distance of $5.5^{+1.4}_{-1.1}$ kpc, this would correspond to a luminosity of $L_{\rm bol}=1.8^{+1.1}_{-0.7}\times 10^{39}\,\rm{erg\,s^{-1}}$. Assuming a nominal BH mass of $10\,M_{\odot}$, $L_{\rm Edd}\approx 1.3\times 10^{39}\,\rm{erg\,s^{-1}}$, meaning an ELF of $L_{\rm bol}/L_{\rm Edd}\approx 1.4^{+0.9}_{-0.6}$ in this case. Black holes with higher predicted natal kicks are expected to be less massive (e.g.; Belczynski & Bulik, 2002; Maccarone et al., 2020), yet reducing the BH mass would only increase this already-large ELF. Therefore, any distance within our $1\sigma$ uncertainties of 4.4 $4\,\text{--}\,6.9$ kpc is set to contravene expectations from Zdziarski et al. (2025) who note that the highest observed luminosities of low-mass XRBs in the hard or hard-intermediate states reside in the regime $L\,\lesssim\,0.3\,L_{\rm Edd}$. Zdziarski et al. (2025) also noted that J1727 underwent a soft-to-hard state transition at very low flux in the period February–April 2024, at $F_{\rm bol}\approx 10^{-8}\,{\rm\,erg\,cm^{-2}\,s^{-1}}$. With this flux, our same distance corresponds to a soft-to-hard state transition ELF of $0.029^{+0.016}_{-0.010}$ assuming a BH mass of $10\,M_{\odot}$, which is more in line with expectations of $1\text{--}4\%$ (Kalemci et al., 2013; Vahdat Motlagh et al., 2019). These ELFs imply that the distance to J1727 is likely on the lower end of our $1\sigma$ range. However, no single distance can satisfy the expectations on both the peak hard state luminosity and soft-to-hard state transition luminosity. While ELFs from soft-to-hard state transitions appear to be steadier (Kalemci et al., 2013), we caution that both the above ELFs are merely indicative and are unlikely to be reliable without a more accurate constraint on the BH mass.
@@ 5 Conclusions
Using H i absorption data from MeerKAT observations of the outburst of Swift J1727.8-1613, we determine a maximum absorption velocity of $24.8\pm 2.8$ km s-1. The higher-velocity absorption seen towards the extragalactic reference source PKS J1733-1304 allows us to use the near kinematic distance as a lower bound of $3.6\pm 0.3\,({stat})\pm 2.3\,({sys})$ kpc, accounting for the systematic uncertainty due to the low Galactic longitude of the source. However, its high Galactic latitude likely implies that we run out of H i clouds along the line of sight.
Making use of a near-UV spectrum of Swift J1727.8-1613 as observed by the Space Telescope Imaging Spectrograph aboard the Hubble Space Telescope, we measure the colour excess or reddening to be $E(B\!-\!V)=0.37\pm 0.01\,({stat})\pm 0.025\,({sys})$. This is significantly lower than previous constraints on E(BV)E(B\!-\!V) from Mata Sánchez et al. (2024, 2025), but it is in good agreement the maximum value derived from Galactic dust maps along this line of sight.
We use E(BV)E(B\!-\!V) to determine the $r$-band extinction, $A_{r}$, and combine this with donor star $r$-band magnitudes provided by Mata Sánchez et al. (2025) for a K4($\pm$1)V main sequence donor star. Under this assumption, we subsequently present $5.5^{+1.4}_{-1.1}$ kpc as the resulting and likely distance to Swift J1727.8-1613, which implies a natal kick velocity of $v_{\rm{kick}}=190\,\pm\,30$ km s-1 and suggests a likely formation in a natal supernova.
If the donor star has instead lost significant mass during the binary evolution, this distance would be smaller. However, the exact evolutionary stage and possible mass loss through accretion of the donor star is unknown. Further observations to better constrain binary inclination angle and the donor star rotational broadening would allow more accurate determination of the primary and secondary masses and consequently the distance.
@@ Acknowledgements
The authors would like to thank James Allison for his development of the ThunderKAT H i Pipeline.
The authors would also like to thank Ilya Mandel, Ryosuke Hirai, Natasha Ivanova, and Thomas Maccarone for useful discussions.
The MeerKAT telescope is operated by the South African Radio Astronomy Observatory, which is a facility of the National Research Foundation, an agency of the Department of Science, Technology and Innovation.
@@
This work made use of the CARTA (Cube Analysis and Rendering Tool for Astronomy) software (Comrie et al. 2024https://cartavis.github.io).
@@
Andrzej Zdziarski acknowledges support from the Polish National Science Center grants 2019/35/B/ST9/03944 and 2023/48/Q/ST9/00138.
Daniel Mata Sánchez acknowledges support via a Ramón y Cajal Fellowship RYC2023-044941.